Puleng receives R1 500,00 from the bank now, that charges 10,5% simple discount per annum. She has to pay back the bank an amount of R1 893,75 in a few months' time. The number of months after which Puleng pays back the loan, rounded to one decimal place, is


1
Expert's answer
2022-03-02T12:55:04-0500

"n=\frac{1-\frac{P}{S}}{i}=\frac{1-\frac{1500}{1893.75}}{0.105}=1.980198"


"1.980198\times12=23.76=23.8"